Solve
dx/x+y-xy^2=dy/x^2-x-y=dz/z(y^2-x^2)
1
Expert's answer
2020-10-09T13:34:21-0400

"\textsf{This is a simultaneous Total}\\\textsf{Differential equation.}\\

\textsf{We solve it by Lagrange's multiplier.}\\

\displaystyle\frac{(l\mathrm{d}x+m\mathrm{d}y +n\mathrm{d}z)}{(l\mathrm{P}+m\mathrm{Q}+n\mathrm{R})} =0.\\

l\mathrm{P}+m\mathrm{Q}+n\mathrm{R} = 0.\hspace{0.1cm}\textsf{So} \hspace{0.1cm}l\mathrm{d}x+m\mathrm{d}y +n\mathrm{d}z =0.\\

\textsf{Here}\hspace{0.1cm} l,m,n \hspace{0.1cm}\textsf{are multipliers.}\\

\textsf{The Lagrange’s auxiliary equations}\\\textsf{for given PDE is}\\

\displaystyle\frac{\mathrm{d}x}{x+y-xy^2}=\frac{\mathrm{d}y}{x^2y-x-y} = \frac{\mathrm{d}z}{z(y^2-x^2)}\\


\displaystyle\textsf{Choosing}\hspace{0.1cm} \left(x,y,\frac{1}{z}\right) \hspace{0.1cm}\\ \textsf{as multipliers, we have}\\

x\mathrm{d}x + y\mathrm{d}y + \frac{1}{z} \mathrm{d}z = 0\\

\textsf{Integrating both sides, we have thus,}\\


\int x\mathrm{d}x + \int y\mathrm{d}y + \int \frac{1}{z} \mathrm{d}z = 0\\

x² + y² + 2\ln(z) = C_1 \\

\textsf{From the first two equations, we have}\\

(x^2y-x-y) \mathrm{d}x - (x+y-xy^2)\mathrm{d}y= 0\\

\textsf{Solving the above DE, we have}\\

(x^2 y - x) \mathrm{d}x - (y - xy^2)\mathrm{d}y - y\mathrm{d}x - x\mathrm{d}y = 0\\

(x^2 y - x) \mathrm{d}x - (y - xy^2)\mathrm{d}y = y\mathrm{d}x + x\mathrm{d}y \\

x(xy - 1) \mathrm{d}x - y(1 - xy)\mathrm{d}y = y\mathrm{d}x + x\mathrm{d}y \\

(xy - 1)(x\mathrm{d}x + y\mathrm{d}y) = y\mathrm{d}x + x\mathrm{d}y \\

\displaystyle x\mathrm{d}x + y\mathrm{d}y = \frac{y\mathrm{d}x + x\mathrm{d}y}{xy - 1}\\

\displaystyle \int x\mathrm{d}x + \int y\mathrm{d}y = \int \frac{\mathrm{d}(xy)}{xy - 1}\\

\displaystyle\frac{x^2}{2} + \frac{y^2}{2} = \log{(xy - 1)} + C_2\\

\displaystyle\frac{x^2}{2} + \frac{y^2}{2} - \log{(xy - 1)} = C_2\\


\textsf{Thus, the solution of the given PDE is}\\
\phi\left(x² + y² + 2\ln(z),\right.\\\left.\frac{x^2}{2} + \frac{y^2}{2} - \log{(xy - 1)}\right) = 0 \hspace{0.2cm}\forall C_1, C_2\in \mathbb{R}"